OpenStudy (anonymous):

help please:) medal rewarded to first best answer. What happens to the volume of a square pyramid if the base dimensions remain the same, but the height is cut to one third of the original? The volume becomes 1/9 of original. The volume becomes 1/3 of original. The volume becomes 3 times larger than original. The volume becomes 9 times larger than original

OpenStudy (mathstudent55):

The volume of a square pyramid is V = (1/3)Ah where A = area of base and h = height For a pyramid with base area A and height h, V = (1/3)Ah If you now substitute h with h/3, you get V = (1/3)(Ah/3) = (1/9)Ah (1/9)Ah is one-third of (1/3)Ah

OpenStudy (anonymous):

thank you!:)

OpenStudy (mathstudent55):

You're welcome